OpenStudy (anonymous):

Lilly takes a train each day to work that averages 35 miles per hour. On her way home, her train ride follows the same path and averages 45 miles per hour. If the total trip takes 2.5 hours, which equation can be used to find n, the number of miles Lilly’s home is from her work?

OpenStudy (shamil98):

I am guessing there are answer choices?

OpenStudy (shamil98):

In this case if it takes 2.5 hours to get there and back, you would find the average 35 + 45 / 2 = 40 mph 40 x 2.5 =

OpenStudy (anonymous):

the options are \[\frac{ n }{ 35} * \frac{ n }{ 45 } = 2.5\] \[\frac{ n }{ 35 } + \frac{ n }{ 45 } = 2.5\] 35n + 45n = 2.5 35n * 45n = 2.5

OpenStudy (shamil98):

sorry , i honestly don't know.. try bumping your question , i'm stumped..

OpenStudy (ranga):

n/35 + n/45 = 2.5

OpenStudy (ranga):

Onward trip is 35 miles in 1 hour. So for n miles it takes n/35 hours. Return trip is 45 miles in 1 hour. So for n miles it takes n/45 hours. Total time for the trip = n/35 + n/45 = 2.5

OpenStudy (anonymous):

thank you